OpenStudy (anonymous):

In early 2012, the price of gas was $3.64. This represents a 40% increase from early 2009. What was the price of gas in early 2009?

OpenStudy (anonymous):

3.64/1.4 = (what you are looking for) How? in 2009 years it represents 100% of whole price, therefore in 2012 it will be 100% + 40% = 140%

OpenStudy (anonymous):

thank you that helped a lot

OpenStudy (anonymous):

... and you make equation : 3.64/1.40 = x/1

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Brian and Ron went to TGI Friday’s for dinner. After adding a 20% tip, they paid a total of $42. What was the cost of the meal before they added the tip?

OpenStudy (anonymous):

yet again, same situation. let's say theyr price of meal is equal 100 %, and you add 20 % extra so : 100% +20% is what you are given. therefore (i'll add example so you get how it's done step by step) x1 - y1 x2 - y2 x1*y2 = x2*y1 and you use that to make: 42 - 120 % x - 100% helps?